Solve the problems and show the complete solution.

At the center of a cube of edge 80.0 cm, a charge of 170 µC is found. (1 µC = 10-6 Coulombs) (Permittivity of free space (εo) 8.85x10-12 C2 / N.m2 )

a) Find the electric flux. (Φ = ∑Qenclosed / E0 )

b) Find the total Φ through each face of the cube. (Φone-face = Φ / 6)

c) Find the Φ through the whole surface of the cube.


1
Expert's answer
2021-02-15T17:39:06-0500

a)


"\Phi=\frac{170\cdot10^{-6}}{8.85\cdot10^{-12}}=1.92\cdot10^{7}\ Wb"

b)


"\frac{1}{6}\Phi=\frac{170\cdot10^{-6}}{6\cdot8.85\cdot10^{-12}}=3.2\cdot10^{6}\ Wb"

c)


"\Phi=\frac{170\cdot10^{-6}}{8.85\cdot10^{-12}}=1.92\cdot10^{7}\ Wb"